OpenStudy (anonymous):

the uniform federal accesability standards specify that a ramp angle use for a wheel chair ramp must be less than or equal to 4.67 degrees. if the length of the ramp is 20 feet and the verticle rise is 17 inches what is the angle going to be

OpenStudy (mathmate):

gradient = tan (theta) = rise/run = 17/(20*12)= 0.0780333 Can you find theta?

OpenStudy (anonymous):

4.052 degrees :)

OpenStudy (anonymous):

=arctan(17/(20*12)), indeed.

OpenStudy (anonymous):

arctan(17/239.4) = 0.0709 in radians or 4.0622708 in degrees